# Sparse juntas on the biased hypercube

### v4

@@ 3.3 Quantized sparse juntas
which shows that $\widetilde{fg}(U)$ is the sum of $3^{|U|}\leq 3^{2d}$ products of two elements from $B$. ∎
The junta agreement theorem will result in a quantized sparse junta $g$ which is close to Boolean in an $L_{0}$ sense: $\Pr[g\notin\{0,1\}]=O(\epsilon)$. In this section, we show how to conclude closeness in an $L_{2}$ sense: $\Pr[\operatorname{dist}(g,\{0,1\})^{2}]=O(\epsilon)$.
@@
For the other direction, Equation 2 implies that $\operatorname{supp}(f)$ is $(d,C/p,\epsilon)$-sparse, where $\epsilon=O_{d,C}(\Pr[f\neq 0])$. Lemma 3.6 shows that $\operatorname{supp}(f^{2})\subseteq\operatorname{supp}(f)^{\sqcup 2}$ is $(2d,K/p,\epsilon)$-sparse, where $K=O_{d,C}(1)$, and Lemma 3.9 shows that $f^{2}$ is $B^{\prime}$-quantized for some finite $B^{\prime}$ depending only on $d,B$. Therefore
$\operatorname*{\mathbb{E}}[f^{2}]\leq\max_{b\in B^{\prime}}b\sum_{e=0}^{2d}(K/p)^{e}\epsilon\cdot p^{e}=O_{d,C,B}(\Pr[f\neq 0]).\qed$
@@
Proof.
Let $g=f(f-1)$. Since $\operatorname{supp}(g)\subseteq\operatorname{supp}(f)^{\sqcup 2}$, Lemmas 3.5 and 3.9 show that $g$ is a $B^{\prime}$-quantized $(2d,K/p)$-sparse junta, where $B^{\prime},K$ depend on $d,C,B$. Therefore Lemma 3.10 implies that
@@
We would like to bound the right-hand side using Lemma 3.7, but it is not necessarily the case that $\mathcal{F}\subseteq\operatorname{supp}(f)$. However, Lemma 2.7 shows that $\mathcal{F}\subseteq\operatorname{supp}(f)^{\sqcup L_{d}}$. Lemma 3.5 shows that $\operatorname{supp}(f)^{\sqcup L_{d}}$ is $(L_{d}d,K^{\prime}/p)$-sparse, where $K^{\prime}$ depends on $d,C$. Therefore Lemma 3.7, applied with $\mathcal{F}$ and $\mathcal{G}=\operatorname{supp}(f)^{\sqcup L_{d}}$, shows that
@@
where $\mathcal{E}$ is the event that $y_{S}=1$ for some $S\in\mathcal{F}$, and $y_{R}=0$ for all $R\in\operatorname{supp}(f)^{\sqcup L_{d}}$ such that $R\not\subseteq S$. If the event $\mathcal{E}$ happens then $f(y)=f(\mathbbm{1}_{S})\notin\{0,1\}$. Moreover, since $|S|\leq L_{d}$ by Lemma 2.7 and $f$ is $B$-quantized, $\operatorname{dist}(f(y),\{0,1\})^{2}=\Omega_{d,B}(1)$. Therefore
$\Pr[f\notin\{0,1\}]=O_{d,C,B}(\operatorname*{\mathbb{E}}[\operatorname{dist}(f,\{0,1\})^{2}]).\qed$
@@ 4 Junta agreement theorem
Our previous work [DFH17] proves a stronger version of Theorem 4.1, which does not require the $g_{S}$ to be juntas but has the same conclusion (without the guarantee that $g$ is a sparse junta, which follows from the $g_{S}$ being juntas). The proof of this stronger version is somewhat non-intuitive, using the notion of good and excellent sets which is borrowed from [IKW12]. It turns out that the proof simplifies dramatically when the $g_{S}$ are juntas, and this is the proof that we present here.
@@
Our first goal is to show that $\Pr[g|_{S}\neq g_{S}]=O_{K,c}(\epsilon)$. We start by applying the union bound to relate the probability to disagreement on a particular monomial coefficient:
@@
We conclude that the number of sets $A\in\operatorname{supp}(g)$ containing $R$ of size $|R|+e$ is at most $2Kq^{-e}\leq(2K/q)^{e}$, as needed. ∎
@@ 5 Structure theorem
Suppose that $f\colon(\{0,1\}^{n},\mu_{p})\to\{0,1\}$ is $\epsilon$-close to degree $d$, where $p\leq 1/2$. Then $\Pr[f\neq g]=O(\epsilon)$ and $\operatorname*{\mathbb{E}}[(f-g)^{2}]=O(\epsilon)$ for some function $g\colon(\{0,1\}^{n},\mu_{p})\to\mathbb{R}$ satisfying the following properties:
@@
    For every $e$, the number of minimal non-Boolean inputs of $g$ of weight $e$ is $O_{d}(\epsilon/p^{e})$.
Conversely, if $g$ satisfies these properties, then it is $O_{d}(\epsilon)$-close to Boolean.
@@
Let $f$ be as in the theorem. Then $f$ is $O_{d}(\epsilon)$-close to $\operatorname{round}(g,\{0,1\})$, where $g$ is the function in the theorem.
Conversely, if $g$ satisfies the conditions in the theorem, then $\operatorname{round}(g,\{0,1\})$ is $O_{d}(\epsilon)$-close to degree $d$.
Proof of Corollary 5.2.
@@
$\Pr[f\neq\operatorname{round}(g,\{0,1\})]\leq\Pr[f\neq g]=O(\epsilon).$
@@
$\operatorname*{\mathbb{E}}[(\operatorname{round}(g,\{0,1\})-g)^{2}]=O_{d}(\epsilon)$
since $g$ is $O_{d}(\epsilon)$-close to Boolean. ∎
@@
    We first construct a sparse junta $h$ which is $O(\epsilon)$-close to $f$.
@@ 5.1 Step 1: Agreement
Since $f$ is $\epsilon$-close to degree $d$, there is a degree $d$ function $f^{\leq d}$ such that $\operatorname*{\mathbb{E}}_{\mu_{p}}[(f-f^{\leq d})^{2}]\leq\epsilon$.
@@
$\epsilon_{S}=\operatorname*{\mathbb{E}}_{\mu_{1/2}}[(f|_{S}-f^{\leq d}|_{S})^{2}],$
@@
$\operatorname*{\mathbb{E}}_{S\sim\mu_{2p}}[\epsilon_{S}]=\operatorname*{\mathbb{E}}_{\mu_{p}}[(f-f^{\leq d})^{2}]\leq\epsilon.$
Applying the Kindler–Safra theorem (Theorem 2.4), there exist Boolean degree $d$ functions $g_{S}$ such that $\operatorname*{\mathbb{E}}_{\mu_{1/2}}[(f|_{S}-g_{S})^{2}]=O(\epsilon_{S})$, and so
$\operatorname*{\mathbb{E}}_{S\sim\mu_{2p}}\left[\operatorname*{\mathbb{E}}_{\mu_{1/2}}[(f|_{S}-g_{S})^{2}]\right]=O\left(\operatorname*{\mathbb{E}}_{S\sim\mu_{2p}}[\epsilon_{S}]\right)=O(\epsilon).$
@@
$\Pr_{(S_{1},S_{2},T)\sim\mu_{2p,\sqrt{2}p}}[g_{S_{1}}|_{T}\neq g_{S_{2}}|_{T}]\leq 2^{K+1}\operatorname*{\mathbb{E}}_{S\sim\mu_{2p}}\left[\operatorname*{\mathbb{E}}_{\mu_{1/2}}[(g_{S}-f|_{S})^{2}]\right]=O(\epsilon),$
@@
To conclude this step of the proof, we need to show that $\operatorname*{\mathbb{E}}_{\mu_{p}}[(f-h)^{2}]=O(\epsilon)$. We first show that $\Pr_{\mu_{p}}[f\neq h]=O(\epsilon)$. Indeed,
$\Pr_{\mu_{p}}[f\neq h]=\operatorname*{\mathbb{E}}_{S\sim\mu_{2p}}\left[\Pr_{\mu_{1/2}}[f|_{S}\neq h|_{S}]\right]\leq\operatorname*{\mathbb{E}}_{S\sim\mu_{2p}}\left[\Pr_{\mu_{1/2}}[f|_{S}\neq g_{S}]+\mathbbm{1}[g_{S}\neq h|_{S}]\right]=O(\epsilon).$
@@
$\operatorname*{\mathbb{E}}_{\mu_{p}}[\operatorname{dist}(h,\{0,1\})^{2}]=O(\Pr_{\mu_{p}}[h\notin\{0,1\}])=O(\Pr_{\mu_{p}}[h\neq f])=O(\epsilon).$
@@
$\operatorname*{\mathbb{E}}_{\mu_{p}}[(f-h)^{2}]\leq 2\operatorname*{\mathbb{E}}_{\mu_{p}}[(f-\operatorname{round}(h,\{0,1\}))^{2}]+2\operatorname*{\mathbb{E}}_{\mu_{p}}[(\operatorname{round}(h,\{0,1\})-h)^{2}]=2\Pr_{\mu_{p}}[f\neq h]+O(\epsilon)=O(\epsilon).$
@@
    $\operatorname*{\mathbb{E}}[(h-f)^{2}]=O(\epsilon)$.
@@ 5.2 Step 2: Fixing coefficients
    $\operatorname*{\mathbb{E}}[(h_{e}-f)^{2}]=O_{d,e}(\epsilon)$.
@@
Lemmas 3.5 and 3.9 show that $h_{e}-h_{e-1}$ is a $B^{+}_{e}$-quantized $(d,O(1/p))$-sparse junta, for some finite set $B^{+}_{e}$. Therefore Lemma 3.10, applied with $f=h_{e}-h_{e-1}$, shows that
$\operatorname*{\mathbb{E}}[(h_{e-1}-h_{e})^{2}]=O(\Pr[h_{e-1}\neq h_{e}])=O(\epsilon).$
@@
$\operatorname*{\mathbb{E}}[(h_{e}-f)^{2}]\leq 2\operatorname*{\mathbb{E}}[(h_{e}-h_{e-1})^{2}]+2\operatorname*{\mathbb{E}}[(h_{e-1}-f)^{2}]=O(\epsilon).$
@@
Taking $g=h_{d}$, the function $g$ satisfies all properties stated in Theorem 5.1, except for Item iii, which states that the number of minimal non-Boolean inputs of $g$ of weight $e$ is $O(\epsilon/p^{e})$.
@@
$\sum_{S\in\mathcal{F}}p^{|S|}=O(\Pr[g\notin\{0,1\}])=O(\Pr[g\neq f])=O(\epsilon).$
Therefore $\mathcal{F}$ contains at most $O(\epsilon/p^{e})$ inputs of weight $e$.
@@ 5.3 Converse part
    $g$ has $O(\epsilon/p^{e})$ minimal non-Boolean inputs of weight $e$, for every $e$.
We need to show that $g$ is $O(\epsilon)$-close to Boolean.
@@
$\Pr[g\notin\{0,1\}]\leq\sum_{e\leq L_{d}}p^{e}\cdot O(\epsilon/p^{e})=O(\epsilon).$
Lemma 3.11 concludes that $g$ is $O(\epsilon)$-close to Boolean.
@@ 6 Monotone structure theorems
Suppose that $f\colon(\{0,1\}^{n},\mu_{p})\to\{0,1\}$ is monotone and $\epsilon$-close to degree $d$, where $p\leq 1/2$. Then $\Pr[f\neq g]=O(\epsilon)$ and $\operatorname*{\mathbb{E}}[(f-g)^{2}]=O(\epsilon)$ for some function $g\colon(\{0,1\}^{n},\mu_{p})\to\mathbb{R}$ satisfying the following properties:
@@
    For every $e$, the number of minimal non-Boolean inputs of $g$ of weight $e$ is $O_{d}(\epsilon/p^{e})$.
@@
    For every $e$, the number of minimal non-monotone Boolean inputs of $g$ of weight $e$ is $O_{d}(\epsilon/p^{e})$.
Conversely, if $g$ satisfies these properties then it is $O_{d}(\epsilon)$-close to some monotone Boolean function.
@@
Let $f$ be as in theorem. Then $f$ is $O_{d}(\epsilon)$-close to $\operatorname{round}(g,\{0,1\})$, where $g$ is the function in the theorem.
Conversely, if $g$ satisfies the conditions in the theorem, then $\operatorname{round}(g,\{0,1\})$ is $O_{d}(\epsilon)$-close to a monotone Boolean function which is $O_{d}(\epsilon)$-close to degree $d$.
Proof of Corollary 6.2.
The first statement follows as in the proof of Corollary 5.2. Now suppose that $g$ satisfies the conditions in the theorem. According to the theorem, $\operatorname*{\mathbb{E}}[(g-f)^{2}]=O_{d}(\epsilon)$ for some monotone Boolean function $f$. Since $g$ has degree $d$, $f$ is $O_{d}(\epsilon)$-close to degree $d$. ∎
The proof of the direct part of Theorem 6.1 closely resembles the proof of the direct part of Theorem 5.1, while the proof of the converse part is a little bit more tricky. For both proofs, we suppress the dependence of big O constants on $d$.
@@
Suppose that $f\colon(\{0,1\}^{n},\mu_{p})\to\{0,1\}$ is monotone and $\epsilon$-close to degree $d$. The argument in Section 5.1 applies, constructing a function $h$ which satisfies the properties stated there (we repeat them below as part of the proof). We now run an argument in the style of Section 5.2, constructing a sequence of functions $h=h_{-1},h_{0},\ldots,h_{d}=g$, where $h_{e}$ satisfies the following properties:
@@
    $\operatorname*{\mathbb{E}}[(h_{e}-f)^{2}]=O_{d,e}(\epsilon)$.
@@
The argument in Section 5.2 shows that the first term on the right is $O(\epsilon)$. In order to bound the second term, we apply the reverse union bound (Lemma 3.7) with $p:=2p$, concluding that
@@
Since the left-hand side is $O(\epsilon)$, we conclude that
$\sum_{S\in\mathcal{F}_{m}}p^{|S|}=O(\Pr_{\mu_{2p}}[\mathcal{E}])=O(\epsilon).$
This shows that $\Pr[h_{e}\neq h_{e-1}]=O(\epsilon)$, and so $\Pr[h_{e}\neq f]\leq\Pr[h_{e}\neq h_{e-1}]+\Pr[h_{e-1}\neq f]=O(\epsilon)$.
@@
    $g$ has $O(\epsilon/p^{e})$ minimal non-Boolean inputs of weight $e$, for all $e$.
@@
    $g$ has $O(\epsilon/p^{e})$ minimal non-monotone Boolean inputs of weight $e$, for all $e$.
We need to show that $g$ is $O(\epsilon)$-close to some monotone Boolean function $f$.
Theorem 5.1 shows that $g$ is $O(\epsilon)$-close to the Boolean function $G=\operatorname{round}(g,\{0,1\})$. However, $G$ is not necessarily monotone. Accordingly, we define
@@
which is monotone by construction. It remains to show that $g$ is $O(\epsilon)$-close to $f$. We will show that $G$ is $O(\epsilon)$-close to $f$. This would imply that
$\operatorname*{\mathbb{E}}[(g-f)^{2}]\leq 2\operatorname*{\mathbb{E}}[(g-G)^{2}]+2\operatorname*{\mathbb{E}}[(G-f)^{2}]=O(\epsilon)+2\Pr[G\neq f]=O(\epsilon).$
@@
$\Pr[G\neq f]\leq\sum_{\mathbbm{1}_{T}\in\mathcal{F}}p^{|T|}\leq\sum_{e\leq\max(L_{d},M_{d})}p^{e}\cdot O(\epsilon/p^{e})=O(\epsilon),$
using Lemmas 2.7 and 2.8 to bound the size of an input in $\mathcal{F}$. This concludes the proof.

### v6

@@ 3.3 Quantized sparse juntas
which shows that $\widetilde{fg}(U)$ is the sum of $3^{|U|}\leq 3^{2d}$ products of two elements from $B$.
The junta agreement theorem will result in a quantized sparse junta $g$ which is close to Boolean in an $L_{0}$ sense: $\Pr[g\notin\{0,1\}]=O(\varepsilon)$. In this section, we show how to conclude closeness in an $L_{2}$ sense: $\Pr[\operatorname{dist}(g,\{0,1\})^{2}]=O(\varepsilon)$.
@@
For the other direction, Equation 2 implies that $\operatorname{supp}(f)$ is $(d,C/p,\varepsilon)$-sparse, where $\varepsilon=O_{d,C}(\Pr[f\neq 0])$. Lemma 3.7 shows that $\operatorname{supp}(f^{2})\subseteq\operatorname{supp}(f)^{\sqcup 2}$ is $(2d,K/p,\varepsilon)$-sparse, where $K=O_{d,C}(1)$, and Lemma 3.12 shows that $f^{2}$ is $B^{\prime}$-quantized for some finite $B^{\prime}$ depending only on $d,B$. Therefore
$\operatorname*{\mathbb{E}}[f^{2}]\leq\max_{b\in B^{\prime}}b\sum_{e=0}^{2d}(K/p)^{e}\varepsilon\cdot p^{e}=O_{d,C,B}(\Pr[f\neq 0]).$
@@
Proof 3.17.
Let $g=f(f-1)$. Since $\operatorname{supp}(g)\subseteq\operatorname{supp}(f)^{\sqcup 2}$, Lemmas 3.5 and 3.12 show that $g$ is a $B^{\prime}$-quantized $(2d,K/p)$-sparse junta, where $B^{\prime},K$ depend on $d,C,B$. Therefore, Lemma 3.14 implies that
@@
We would like to bound the right-hand side using Lemma 3.9, but it is not necessarily the case that $\mathcal{F}\subseteq\operatorname{supp}(f)$. However, Lemma 2.11 shows that $\mathcal{F}\subseteq\operatorname{supp}(f)^{\sqcup L_{d}}$. Lemma 3.5 shows that $\operatorname{supp}(f)^{\sqcup L_{d}}$ is $(L_{d}d,K^{\prime}/p)$-sparse, where $K^{\prime}$ depends on $d,C$. Therefore, Lemma 3.9, applied with $\mathcal{F}$ and $\mathcal{G}=\operatorname{supp}(f)^{\sqcup L_{d}}$, shows that
@@
where $\mathcal{E}$ is the event that $y_{S}=1$ for some $S\in\mathcal{F}$, and $y_{R}=0$ for all $R\in\operatorname{supp}(f)^{\sqcup L_{d}}$ such that $R\not\subseteq S$. If the event $\mathcal{E}$ happens then $f(y)=f(\mathbbm{1}_{S})\notin\{0,1\}$. Moreover, since $|S|\leq L_{d}$ by Lemma 2.11 and $f$ is $B$-quantized, $\operatorname{dist}(f(y),\{0,1\})^{2}=\Omega_{d,B}(1)$. Therefore,
$\Pr[f\notin\{0,1\}]=O_{d,C,B}(\operatorname*{\mathbb{E}}[\operatorname{dist}(f,\{0,1\})^{2}]).$
@@ 4 Junta agreement theorem
Our previous work [DinurFH-agree] proves a stronger version of Theorem 4.1, which does not require the $g_{S}$ to be juntas but has the same conclusion (without the guarantee that $g$ is a sparse junta, which follows from the $g_{S}$ being juntas). The proof of this stronger version is somewhat non-intuitive, using the notion of good and excellent sets which is borrowed from [ImpagliazzoKW2012]. It turns out that the proof simplifies dramatically when the $g_{S}$ are juntas, and this is the proof that we present here.
@@
Our first goal is to show that $\Pr[g|_{S}\neq g_{S}]=O_{K,c}(\varepsilon)$. We start by applying the union bound to relate the probability to disagreement on a particular monomial coefficient:
@@
We conclude that the number of sets $A\in\operatorname{supp}(g)$ containing $R$ of size $|R|+e$ is at most $2Kq^{-e}\leq(2K/q)^{e}$, as needed.
@@ 5 Structure theorem
Suppose that $f\colon(\{0,1\}^{n},\mu_{p})\to\{0,1\}$ is $\varepsilon$-close to degree $d$, where $p\leq 1/2$. Then $\Pr[f\neq g]=O(\varepsilon)$ and $\operatorname*{\mathbb{E}}[(f-g)^{2}]=O(\varepsilon)$ for some function $g\colon(\{0,1\}^{n},\mu_{p})\to\mathbb{R}$ satisfying the following properties:
@@
    For every $e$, the number of minimal non-Boolean inputs of $g$ of weight $e$ is $O_{d}(\varepsilon/p^{e})$.
Conversely, if $g$ satisfies these properties, then it is $O_{d}(\varepsilon)$-close to Boolean.
@@
Let $f$ be as in the theorem. Then $f$ is $O_{d}(\varepsilon)$-close to $\operatorname{round}(g,\{0,1\})$, where $g$ is the function in the theorem.
Conversely, if $g$ satisfies the conditions in the theorem, then $\operatorname{round}(g,\{0,1\})$ is $O_{d}(\varepsilon)$-close to degree $d$.
Proof 5.3 (Proof of Corollary 5.2).
@@
$\Pr[f\neq\operatorname{round}(g,\{0,1\})]\leq\Pr[f\neq g]=O(\varepsilon).$
@@
$\operatorname*{\mathbb{E}}[(\operatorname{round}(g,\{0,1\})-g)^{2}]=O_{d}(\varepsilon)$
since $g$ is $O_{d}(\varepsilon)$-close to Boolean.
@@
    We first construct a sparse junta $h$ which is $O(\varepsilon)$-close to $f$.
@@ 5.1 Step 1: Agreement
Since $f$ is $\varepsilon$-close to degree $d$, there is a degree $d$ function $f^{\leq d}$ such that $\operatorname*{\mathbb{E}}_{\mu_{p}}[(f-f^{\leq d})^{2}]\leq\varepsilon$.
@@
$\varepsilon_{S}=\operatorname*{\mathbb{E}}_{\mu_{1/2}}[(f|_{S}-f^{\leq d}|_{S})^{2}],$
@@
$\operatorname*{\mathbb{E}}_{S\sim\mu_{2p}}[\varepsilon_{S}]=\operatorname*{\mathbb{E}}_{\mu_{p}}[(f-f^{\leq d})^{2}]\leq\varepsilon.$
Applying the Kindler–Safra theorem (Theorem 2.5), there exist Boolean degree $d$ functions $g_{S}$ such that $\operatorname*{\mathbb{E}}_{\mu_{1/2}}[(f|_{S}-g_{S})^{2}]=O(\varepsilon_{S})$, and so
$\operatorname*{\mathbb{E}}_{S\sim\mu_{2p}}\left[\operatorname*{\mathbb{E}}_{\mu_{1/2}}[(f|_{S}-g_{S})^{2}]\right]=O\left(\operatorname*{\mathbb{E}}_{S\sim\mu_{2p}}[\varepsilon_{S}]\right)=O(\varepsilon).$
@@
$\Pr_{(S_{1},S_{2},T)\sim\mu_{2p,\sqrt{2}p}}[g_{S_{1}}|_{T}\neq g_{S_{2}}|_{T}]\leq 2^{K+1}\operatorname*{\mathbb{E}}_{S\sim\mu_{2p}}\left[\operatorname*{\mathbb{E}}_{\mu_{1/2}}[(g_{S}-f|_{S})^{2}]\right]=O(\varepsilon),$
@@
To conclude this step of the proof, we need to show that $\operatorname*{\mathbb{E}}_{\mu_{p}}[(f-h)^{2}]=O(\varepsilon)$. We first show that $\Pr_{\mu_{p}}[f\neq h]=O(\varepsilon)$. Indeed,
$\Pr_{\mu_{p}}[f\neq h]=\operatorname*{\mathbb{E}}_{S\sim\mu_{2p}}\left[\Pr_{\mu_{1/2}}[f|_{S}\neq h|_{S}]\right]\leq\operatorname*{\mathbb{E}}_{S\sim\mu_{2p}}\left[\Pr_{\mu_{1/2}}[f|_{S}\neq g_{S}]+\mathbbm{1}[g_{S}\neq h|_{S}]\right]=O(\varepsilon).$
@@
$\operatorname*{\mathbb{E}}_{\mu_{p}}[\operatorname{dist}(h,\{0,1\})^{2}]=O(\Pr_{\mu_{p}}[h\notin\{0,1\}])=O(\Pr_{\mu_{p}}[h\neq f])=O(\varepsilon).$
@@
$\operatorname*{\mathbb{E}}_{\mu_{p}}[(f-h)^{2}]\leq 2\operatorname*{\mathbb{E}}_{\mu_{p}}[(f-\operatorname{round}(h,\{0,1\}))^{2}]+2\operatorname*{\mathbb{E}}_{\mu_{p}}[(\operatorname{round}(h,\{0,1\})-h)^{2}]=2\Pr_{\mu_{p}}[f\neq h]+O(\varepsilon)=O(\varepsilon).$
@@
    $\operatorname*{\mathbb{E}}[(h-f)^{2}]=O(\varepsilon)$.
@@ 5.2 Step 2: Fixing coefficients
    $\operatorname*{\mathbb{E}}[(h_{e}-f)^{2}]=O_{d,e}(\varepsilon)$.
@@
Lemmas 3.5 and 3.12 show that $h_{e}-h_{e-1}$ is a $B^{+}_{e}$-quantized $(d,O(1/p))$-sparse junta, for some finite set $B^{+}_{e}$. Therefore Lemma 3.14, applied with $f=h_{e}-h_{e-1}$, shows that
$\operatorname*{\mathbb{E}}[(h_{e-1}-h_{e})^{2}]=O(\Pr[h_{e-1}\neq h_{e}])=O(\varepsilon).$
@@
$\operatorname*{\mathbb{E}}[(h_{e}-f)^{2}]\leq 2\operatorname*{\mathbb{E}}[(h_{e}-h_{e-1})^{2}]+2\operatorname*{\mathbb{E}}[(h_{e-1}-f)^{2}]=O(\varepsilon).$
@@
Taking $g=h_{d}$, the function $g$ satisfies all properties stated in Theorem 5.1, except for Item iii, which states that the number of minimal non-Boolean inputs of $g$ of weight $e$ is $O(\varepsilon/p^{e})$.
@@
$\sum_{S\in\mathcal{F}}p^{|S|}=O(\Pr[g\notin\{0,1\}])=O(\Pr[g\neq f])=O(\varepsilon).$
Therefore $\mathcal{F}$ contains at most $O(\varepsilon/p^{e})$ inputs of weight $e$.
@@ 5.3 Converse part
    $g$ has $O(\varepsilon/p^{e})$ minimal non-Boolean inputs of weight $e$, for every $e$.
We need to show that $g$ is $O(\varepsilon)$-close to Boolean.
@@
$\Pr[g\notin\{0,1\}]\leq\sum_{e\leq L_{d}}p^{e}\cdot O(\varepsilon/p^{e})=O(\varepsilon).$
Lemma 3.16 concludes that $g$ is $O(\varepsilon)$-close to Boolean.
@@ 6 Monotone structure theorems
Suppose that $f\colon(\{0,1\}^{n},\mu_{p})\to\{0,1\}$ is monotone and $\varepsilon$-close to degree $d$, where $p\leq 1/2$. Then $\Pr[f\neq g]=O(\varepsilon)$ and $\operatorname*{\mathbb{E}}[(f-g)^{2}]=O(\varepsilon)$ for some function $g\colon(\{0,1\}^{n},\mu_{p})\to\mathbb{R}$ satisfying the following properties:
@@
    For every $e$, the number of minimal non-Boolean inputs of $g$ of weight $e$ is $O_{d}(\varepsilon/p^{e})$.
@@
    For every $e$, the number of minimal non-monotone Boolean inputs of $g$ of weight $e$ is $O_{d}(\varepsilon/p^{e})$.
Conversely, if $g$ satisfies these properties then it is $O_{d}(\varepsilon)$-close to some monotone Boolean function.
@@
Let $f$ be as in theorem. Then $f$ is $O_{d}(\varepsilon)$-close to $\operatorname{round}(g,\{0,1\})$, where $g$ is the function in the theorem.
Conversely, if $g$ satisfies the conditions in the theorem, then $\operatorname{round}(g,\{0,1\})$ is $O_{d}(\varepsilon)$-close to a monotone Boolean function which is $O_{d}(\varepsilon)$-close to degree $d$.
Proof 6.3 (Proof of Corollary 6.2).
The first statement follows as in the proof of Corollary 5.2. Now suppose that $g$ satisfies the conditions in the theorem. According to the theorem, there exists some monotone Boolean function $f$ such that $\operatorname*{\mathbb{E}}[(g-f)^{2}]=O_{d}(\varepsilon)$. Since $g$ has degree $d$, $f$ is $O_{d}(\varepsilon)$-close to degree $d$.
The proof of the direct part of Theorem 6.1 closely resembles the proof of the direct part of Theorem 5.1, while the proof of the converse part is slightly more tricky. For both proofs, we suppress the dependence of big $O$constants on $d$.
@@
Suppose that $f\colon(\{0,1\}^{n},\mu_{p})\to\{0,1\}$ is monotone and $\varepsilon$-close to degree $d$. The argument in Section 5.1 applies, constructing a function $h$ which satisfies the properties stated there (we repeat them below as part of the proof). We now run an argument in the style of Section 5.2, constructing a sequence of functions $h=h_{-1},h_{0},\ldots,h_{d}=g$, where $h_{e}$ satisfies the following properties:
@@
    $\operatorname*{\mathbb{E}}[(h_{e}-f)^{2}]=O_{d,e}(\varepsilon)$.
@@
The argument in Section 5.2 shows that the first term on the right is $O(\varepsilon)$. In order to bound the second term, we apply the reverse union bound (Lemma 3.9) with $p:=2p$, concluding that
@@
Since the left-hand side is $O(\varepsilon)$, we conclude that
$\sum_{S\in\mathcal{F}_{m}}p^{|S|}=O(\Pr_{\mu_{2p}}[\mathcal{E}])=O(\varepsilon).$
This shows that $\Pr[h_{e}\neq h_{e-1}]=O(\varepsilon)$, and so $\Pr[h_{e}\neq f]\leq\Pr[h_{e}\neq h_{e-1}]+\Pr[h_{e-1}\neq f]=O(\varepsilon)$.
@@
    $g$ has $O(\varepsilon/p^{e})$ minimal non-Boolean inputs of weight $e$, for all $e$.
@@
    $g$ has $O(\varepsilon/p^{e})$ minimal non-monotone Boolean inputs of weight $e$, for all $e$.
We need to show that $g$ is $O(\varepsilon)$-close to some monotone Boolean function $f$.
Theorem 5.1 shows that $g$ is $O(\varepsilon)$-close to the Boolean function $G=\operatorname{round}(g,\{0,1\})$. However, $G$ is not necessarily monotone. Accordingly, we define
@@
which is monotone by construction. It remains to show that $g$ is $O(\varepsilon)$-close to $f$. We will show that $G$ is $O(\varepsilon)$-close to $f$. This would imply that
$\operatorname*{\mathbb{E}}[(g-f)^{2}]\leq 2\operatorname*{\mathbb{E}}[(g-G)^{2}]+2\operatorname*{\mathbb{E}}[(G-f)^{2}]=O(\varepsilon)+2\Pr[G\neq f]=O(\varepsilon).$
@@
$\Pr[G\neq f]\leq\sum_{\mathbbm{1}_{T}\in\mathcal{F}}p^{|T|}\leq\sum_{e\leq\max(L_{d},M_{d})}p^{e}\cdot O(\varepsilon/p^{e})=O(\varepsilon),$
using Lemmas 2.11 and 2.13 to bound the size of an input in $\mathcal{F}$. This concludes the proof.